OpenStudy (anonymous):

Ideal Body Weight The ideal body weight W for men (in kilograms) as a function of height h(in inches) is given by the function W(h) = 50 + 2.3(h – 60) (a) What is the ideal weight of a 6-foot male? (b) Express the height h as a function of weight W. (c) What is the height of a male who is at his ideal weight of 80 kilograms?

OpenStudy (anonymous):

Hi, a) In order to find the weight, we can substitute the corresponding height (6ft = 12*6 inches= 72 inches) into the equation. So from that we get: W = 50 + 2.3(72 - 60) = 77.6kg b) The question is basically asking us to rearrange the equation and solve for h: W(h) = 50 + 2.3(h - 60) W(h) - 50 = 2.3(h - 60) \[\frac{ W(h) - 50 }{ 2.3 }=h-60\] \[h=\frac{ W(h) - 50 }{ 2.3 }+60\] c) We can use the equation from b and substitute the ideal weight ( W(h) = 80kg) that would give us: \[h=\frac{ 80 - 50 }{ 2.3 }+60\] h = 73 inches = 6.1 ft Hopefully that makes sense!